### v1

# SplatTalk: 3D VQA with Gaussian Splatting

Abstract
Language-guided 3D scene understanding is important for advancing applications in robotics, AR/VR, and human-computer interaction, enabling models to comprehend and interact with 3D environments through natural language. While 2D vision-language models (VLMs) have achieved remarkable success in 2D VQA tasks, progress in the 3D domain has been significantly slower due to the complexity of 3D data and the high cost of manual annotations. In this work, we introduce SplatTalk, a novel method that uses a generalizable 3D Gaussian Splatting (3DGS) framework to produce 3D tokens suitable for direct input into a pretrained LLM, enabling effective zero-shot 3D visual question answering (3D VQA) for scenes with only posed images. During experiments on multiple benchmarks, our approach outperforms both 3D models trained specifically for the task and previous 2D-LMM-based models utilizing only images (our setting), while achieving competitive performance with state-of-the-art 3D LMMs that additionally utilize 3D inputs.
1 Introduction
The rise of Large Multimodal Models (LMMs), following the success of Large Language Models (LLMs), has become increasingly significant in a world where visual understanding is crucial. While 2D LMMs have seen rapid improvements, enabling models to analyze images and answer complex questions [43, 30, 23, 40], progress in 3D LMMs has been slower. This gap is largely due to the scarcity of 3D data compared to 2D datasets. Further, the availability of language-annotated 3D data is even more limited, as annotating 3D scenes since it is significantly more complex and expensive than labeling 2D images. These challenges have hindered the development of language-guided 3D reasoning, making 3D LMMs an important yet underexplored area in multimodal research. In this work, we focus on 3D Visual Question Answering (3D VQA) as it provides a strong benchmark for evaluating a model’s ability to understand and reason about 3D scenes.
Recently, significant progress has been made in 3D multimodal learning, particularly for indoor scenes [19, 16, 29, 52]. Several approaches rely on posed images and 3D inputs, like meshes [44], point clouds [9, 19, 37] and voxels [14], to associate 3D geometry with language features. As such, they have limited applicability in scenarios where a full 3D surface reconstruction is not available.
Previous works have explored integrating semantic features into NeRF [21] or 3D Gaussian Splatting (3DGS) [38], which rely only upon RGB image inputs. However, research on effectively associating 3D geometry with language fetures remains limited. Some methods align individual 3D locations with language features via implicit functions optimized individually for every scene [21, 10]. Others employ SAM [22] to first segment objects and object parts, followed by CLIP [39] to extract object-centric features from the generated masks. These approaches allows models to query information about individual points and/or objects. However, it disregards spatial relationships between objects, making it unsuitable for 3D VQA spatial reasoning tasks.
More recent work [23] has demonstrated the power of LLMs to answer VQA questions using only a sampling of tokens extracted from images, without any underlying 3D representation. While this approach is inspiring, it relies upon the LLM to make sense of a set of tokens that are arranged according to images rather than elements in the 3D scene. As a result, it struggles to answer queries about specific 3D locations (e.g., “What is here?”) and questions requiring a holistic understanding of the entire 3D scene, where the answer depends on establishing correspondences between objects across multiple images (e.g., “What is opposite the door at the other end of the room”?). In this work, we aim to address these issues by consolidating tokens in 3D before providing them to the LLM.
Our approach is called SplatTalk. Given a set of posed RGB images, we integrate language features extracted from the RGB images to build a 3D Gaussian Splatting representation, from which 3D tokens can be extracted for direct input into a pretrained LLM to answer challenging 3D VQA queries (see Fig. 1). The main research contributions are the new methods for 1) integrating language features into the Gaussian representation, and 2) extracting tokens that can be provided directly to an LLM. These methods avoid the issues of concurrent work, ChatSplat [6], which produces tokens that are not directly compatible with an LLM, making language-based reasoning more difficult.
Overall, SplatTalk has a unique combination of three advantages. First, it does not require any 3D surface representation as input. Second, it leverages the full power of a pre-trained LLM to answer difficult 3D VQA questions. Third, it produces a concise 3D representation that can be queried and/or edited in 3D (e.g., cropped to a 3D region, rendered from a novel view, etc.). During experiments with 3D VQA benchmarks, we find that it outperforms 3D models trained only for the benchmark tasks, demonstrating the value of utilizing pre-trained LLMs. SplatTalk further outperforms prior image-only methods using LLMs, showing the value of building a 3D scene representation. Finally, it even achieves competitive performance with SOTA 3D LMMs that utilize 3D scene representation inputs.
In summary, our key contributions are:
We introduce the first self-supervised 3D Gaussian-based method for large-scale zero-shot 3D VQA, SplatTalk, requiring only multi-view images without explicit depth, point clouds or 3D-language annotated supervision.
We show that mean feature extraction from 3D Gaussians encodes holistic scene concepts for 3D VQA, supported by theoretical justification.
We introduce entropy-adaptive token sampling, leading to better 3D VQA performance without additional training.
We demonstrate that SplatTalk outperforms 2D LMMs and achieves performance on par with 3D models using point clouds on various 3D VQA benchmarks, demonstrating the effectiveness of our 3D-aware representations.
2 Related Work
2D LMMs. With advancements in large language models (LLMs), the ability of models to comprehend images and respond to language prompts related to their content has improved significantly. Notably, SOTA models [35, 12, 1, 33, 23] have demonstrated remarkable capabilities in processing multiple images of a 3D scene and answering questions about the shared content across these images. Although multiple RGB images inherently contain 3D information, they do not explicitly model 3D understanding, potentially leading to sub-optimal performance on tasks that require explicit 3D reasoning. In this work, we build upon the strong performance of 2D vision-language models (2D-VLMs), particularly LLaVA-OneVision [23], and integrate them with the 3D priors provided by the 3DGS framework to improve 3D spatial understanding capabilities.
3D LMMs. Recent advancements in 3D indoor scene understanding have explored integrating 3D priors into LLMs using inputs such as 3D point clouds [19, 7] or multi-view RGB images combined with depth or 3D point clouds [52]. These approaches employ various strategies to embed 3D structural information and language features into a shared representation space. While most methods [52, 19] fine-tune pretrained visual encoders (both 2D and 3D) alongside LoRA parameters or even the full set of LLM parameters, others [26, 20] opt for a more lightweight approach, training only a projection layer to map visual features into the language feature space. In contrast, our method does not require any explicit depth or 3D inputs. Instead, we leverage the 3DGS framework, which has shown strong capabilities in representing 3D scenes using only multi-view RGB images.
Generalizable 3D Gaussian Splatting. This body of work explores generalizable 3DGS as an alternative to the per-scene optimization paradigm used in traditional 3DGS methods, aiming to improve novel view synthesis across unseen scenes. FreeSplat [46] introduces a cost-volume-based approach to predict 3D Gaussian parameters from sparse-view inputs, improving reconstruction quality without requiring per-scene optimization. One of the key contributions of this work is the introduction of 3D Gaussian triplets, which are latent feature vectors that serve as a basis representation of the scene, capturing the scene information in a compact form. The method then employs a Pixel-wise Triplet Fusion module, which leverages a lightweight GRU network to fuse overlapping 3D Gaussian triplets from multiple views. This process significantly reduces 3D Gaussian redundancy.
Similarly, [8, 5] leverage multi-view stereo (MVS) cues to refine depth estimation and improve geometric consistency in 3D Gaussian representations. These methods primarily focus on RGB novel view synthesis from sparse inputs (typically views). Our work builds on the SOTA model, FreeSplat [46], but diverges significantly in both objectives and methodology. First, while previous methods focus on photorealistic RGB reconstruction, we aim at semantic 3D scene understanding. Second, instead of reconstructing only a localized portion of the scene as shown in these works, our goal is to represent entire 3D environments, capturing both spatial and semantic relationships holistically.
Language-guided 3D Gaussian Splatting. LangSplat [38] is one of the first approaches that integrates semantic features into 3DGS. However, this method requires training on individual scenes, which is significantly expensive, and lacks generalization to unseen scenes. Similarly, follow-up works [48, 25] explore 3DGS for open-vocabulary object segmentation but do not focus on reasoning tasks like 3D VQA. Further, they primarily rely on scene-specific training, limiting their scalability. In contrast, we propose a fully self-supervised framework that learns 3D-aware visual representations without requiring per-scene training. While some works [45, 13] explores a generalizable 3DGS and semantic modeling approach, they primarily focus on open-vocabulary segmentation, without addressing the broader challenges of free-form language reasoning required for 3D VQA.
ChatSplat [6] is a concurrent work that explores natural language conversation within the 3DGS framework. However, it optimizes on a per-scene basis and requires rendering 3D features into 2D feature maps during inference. Additionally, its training pipeline is not fully end-to-end, as it first trains the RGB Gaussians separately before incorporating language features. In contrast, our approach leverages scene latent features that encapsulate both RGB and semantic information, enabling the joint optimization of both modalities in a fully end-to-end manner.
3 Method
The problem we address is formulated as follows: Given a sequence of posed RGB images captured from a 3D indoor scene, along with language prompts related to the scene’s content, the goal is to generate the corresponding language responses. That is, we address 3D VQA tasks with solely posed RGB images as inputs.
We provide an overview of our approach in Fig. 2, which consists of three main steps: 1) Training the feature autoencoder, 2) Self-supervised 3D-Language Gaussian Splatting Training and 3) 3D VQA Inference. First, we extract the high-dimensional language features from LLaVA-OV [23] for each RGB image by using the pretrained image encoder and multimodal projector. Since these features are high-dimensional, unbounded and sparse, we train an autoconder to project them into a more compact, low-dimensional space. These low-dimensional features then serve as pseudo ground-truth for training the language representations in SplatTalk. SplatTalk architecture is a feed-forward network consists of a Gaussian encoder and decoder that predicts Gaussian parameters for rendering. During 3D VQA inference stage, we extract the trained language features from the 3D Gaussians and directly feed them into the LLM, along with the language prompt, to generate free-form responses.
3.1 Integrating Language Features
An important step of our computational pipeline is the self-supervised method to integrate language features into a generalizable 3DGS framework. We identify four key factors that shape our approach for this step: (1) Leveraging features that align with the LLM’s input space for direct language reasoning, (2) Efficient feature dimensionality reduction for effective CUDA-based rendering, and (3) Joint training of RGB and semantic features for coherent multimodal representation learning.
Visual Tokens for 2D Pseudo Ground-truth Features. We extract the visual tokens used as inputs to the LLM for training our 3DGS framework. Our empirical findings indicate that extracting features directly from the visual encoder for training 3DGS requires retraining the multimodal projector layer which aligns the visual with language spaces. This observation is further corroborated by [6], which demonstrates that features extracted before the multimodal projector cannot be seamlessly integrated into the language model. A possible explanation is that the image encoder’s native feature space (before the projector) is high-dimensional and unbounded, leading to sparser features with greater variance across different dimensions. As a result, the projector becomes more sensitive to subtle variations in the input features, making it difficult to directly map them into a token representation that aligns with the LLM’s understanding. In contrast, the visual tokens generated after the projector are already well-aligned with the language model’s latent space, allowing the LLM to interpret and reason with them directly.
Feature Dimensionality Reduction. Building on the previous factor, directly optimizing these language embeddings within a differentiable rendering pipeline is challenging because they are extremely high-dimensional (e.g., dimensions), sparse, and contain unbounded values, making learning both unstable and impractical. Therefore, we learn a generalizable autoencoder that projects these sparse, high-dimensional visual tokens into a compact unit hypersphere with much lower dimension (e.g., dimensions), ensuring feature smoothness. Note that the language representations from LLaVA, which support reasoning and free-form language comprehension, are significantly more complex than those from CLIP or segmentation-based models like SAM. Different from prior works that compress semantic features into extremely low-dimensional vectors (e.g., dimensions [38, 48]), which simplifies rendering but results in significant information loss, our approach prioritizes preserving rich language features by maintaining a higher-dimensional representation (e.g., ). To efficiently render high-dimensional feature maps on CUDA, we draw inspiration from [51] and implement a parallel CUDA rasterizer pipeline, where semantic features and RGB are rendered simultaneously, using shared Gaussian parameters.
Unlike previous methods [38, 6] that train a separate autoencoder for each individual scene, we train a single autoencoder across all training scenes, each consisting of multiple frames. This approach improves the autoencoder’s generalizability, reducing computational overhead by eliminating the need for retraining to compress features from unseen scenes. Training the autoencoder separately from the 3DGS model ensures that it is dedicated solely to compressing high-dimensional feature vectors while preserving the essential scene information without being influenced by the complexities of the rendering pipeline. Our empirical findings further indicate that jointly training a feature decoder to project low-dimensional embeddings back to the high-dimensional space introduces significant instabilities in the training process.
Joint-training RGB and Language. In contrast to previous approaches that first train RGB Gaussians and then freeze them before integrating language features [6, 38], we adopt a joint training strategy, where RGB and language features are learned simultaneously. This is due to the fact that visual RGB and semantic features play a complementary role in effectively representing 3D scenes [36]. In feedforward models like FreeSplat, where the scene is first encoded into a Gaussian latent space before being decoded into specific Gaussian parameters for RGB rendering, these Gaussian latent features inherently capture holistic scene information, including both appearance (RGB) and semantics. Based on this observation, we introduce a new Gaussian parameter decoder head that predicts semantic features for each 3D Gaussian alongside rendering parameters, enabling joint optimization. Unlike approaches that explicitly model these interactions [36, 42] using extra cross-attention or projection layers, SplatTalk implicitly integrates semantic and visual information within the Gaussian representation. This paradigm naturally facilitates multimodal interactions between RGB and language features.
To incorporate more views and comprehensively cover the scene, we reduce the input resolution of the RGB images. This adjustment aligns well with the inherently low spatial resolution of the visual tokens extracted from LLaVA-OV, and also prevents the model from disproportionately allocating resources to modeling RGB features, thus encouraging a balanced representation of semantic and RGB information.
| Method | Modality | ScanQA (val) | SQA3D (test) | |||||
| CIDEr | METEOR | ROUGE | EM@1 | EM@1-R | EM@1 | EM@1-R | ||
| Specialist Models | ||||||||
| ScanQA [2] | PC | 64.9 | 13.1 | 33.1 | 21.1 | - | 47.2 | - |
| ScanRefer+MCan [49] | PC | 55.4 | 11.5 | 30.0 | 18.6 | - | - | - |
| 3D-VisTA [53] | PC | 69.6 | 13.9 | 35.7 | 22.4 | - | 48.5 | - |
| SQA3D [34] | PC | - | - | - | - | - | 46.6 | - |
| Generalist 3D LMMs | ||||||||
| Chat-Scene [18] | PC+I | 87.7 | 18.0 | 41.6 | 21.6 | - | 54.6 | 57.5 |
| Scene-LLM [14] | PC+I | 80.0 | 15.8 | 35.9 | - | - | 53.6 | - |
| Grounded 3D-LLM [9] | PC | 72.7 | - | - | - | - | - | - |
| Chat-3D-v2 [47] | PC | 77.1 | 16.1 | 40.1 | - | - | 54.7 | - |
| Finetuned 3D LMMs | ||||||||
| 3D-LLM [16] | PC+I | 69.4 | 14.5 | 35.7 | 20.5 | - | - | - |
| LEO [19] | PC+I | 101.4 | 20.0 | 49.2 | 24.5 | 47.6 | 50.0 | 52.4 |
| Scene-LLM [14] | PC+I | 80.0 | 16.6 | 40.0 | 27.2 | - | 54.2 | - |
| LL3DA [7] | PC | 76.8 | 15.9 | 37.3 | - | - | - | - |
| Chat-3D [47] | PC | 53.2 | 11.9 | 28.5 | - | - | - | - |
| 2D LMM-Based | ||||||||
| VideoChat2 [27] | I | 49.2 | 9.5 | 28.2 | 19.2 | - | 37.3 | - |
| GPT-4V [35] | I | 59.6 | 13.5 | 33.4 | - | - | - | - |
| Claude [1] | I | 57.7 | 10.0 | 29.3 | - | - | - | - |
| LLaVA-NeXT-Video [24] | I | 46.2 | 9.1 | 27.8 | - | - | 34.2 | - |
| LLaVA-OV [23] | I | 50.0 | 13.0 | 29.4 | 15.6 | 32.5 | 25.4 | 35.8 |
| SplatTalk | I | 61.7 | 14.2 | 32.7 | 17.1 | 32.2 | 26.1 | 41.3 |
| SplatTalk-ScanQA-FT | I | 77.1 | 15.4 | 38.7 | 22.3 | 38.3 | 42.0 | 44.7 |
| SplatTalk-3DVQA-FT | I | 77.5 | 15.6 | 38.5 | 22.4 | 38.3 | 47.6 | 49.4 |
These design choices establish our approach as a fundamentally novel method for integrating language features into 3DGS, grounded in first-principles observations rather than heuristic adaptations. By addressing the core challenges of multimodal alignment, feature compression, and joint optimization, our approach is not merely a trivial extension of existing works, but a carefully designed framework that ensures both effectiveness and stability in learning rich, spatially-aware semantic representations.
3.2 Extracting 3D Language Features
During inference, we directly extract language features at the mean position of each 3D Gaussian. Although the rendering process weights features according to the Gaussians’ covariances and opacities, we demonstrate that the semantic feature derived from each Gaussian’s mean effectively captures the scene’s semantics. We further validate this choice by highlighting the conceptual similarity between our training optimization process and the Expectation-Maximization (EM) algorithm. For simplification, let be the loss, where is the predicted rendered feature map of frame at pixel and is the pseudo ground-truth feature map obtained from LLaVA-OV. Replacing by the rendering equation and solving for the optimal :
where . Note that this iterative optimization procedure, rendering and minimizing reconstruction errors, is conceptually analogous to the EM algorithm. Specifically, the procedure consists of: (1) an E-step, which estimates each Gaussian’s contribution to pixels across viewpoints given current Gaussian parameters, and (2) an M-step, which updates the Gaussian parameters to minimize reconstruction errors. Consequently, the optimal 3D feature for each Gaussian can be seen as a weighted sum of the 2D feature maps, with the weights representing each Gaussian’s contributions to the pixels, as computed through the rendering process. Therefore the set of 3D Gaussian features is expected to effectively encode the semantic content of the scene as whole from 2D inputs. The extracted 3D language features are then projected back into the original language embedding space using the pretrained decoder and directly provided as inputs to the LLM for downstream 3D-language tasks.
Note that in open-vocabulary segmentation task, querying features at Gaussian centers appear straightforward [4, 38, 48] because the task inherently assumes that each Gaussian corresponds to a distinct, semantically meaningful object or region. These approaches leverage pre-trained SAM-masked CLIP or LSeg embeddings, where features are object-centric, making it possible to retrieve objects simply by computing cosine similarity between a Gaussian’s feature and a text query. However, this assumption does not hold in 3D reasoning tasks. In our case, a single Gaussian does not necessarily encode an explicit object category. Instead, Gaussians contribute to a scene-level representation, where meaning can emerge from the spatial and relational structure of multiple features rather than a direct one-to-one mapping with objects. This makes querying individual Gaussians insufficient for tasks that require context-dependent inference rather than segmentation-based retrieval. Therefore, while our method queries at the mean Gaussians, it is not inherently intuitive or trivial to do so. Our proof using a theoretical connection to the EM algorithm in this section has provided a principled justification for this approach.
| Model | Input | Scene | Setting | Counting | Existence | Attributes | Spatial | Navigation | Others | Overall |
| LEO [19] | T | PC | zero-shot | 0.80 | 15.5 | 11.8 | 7.3 | 2.3 | 15.3 | 7.8 |
| LLaVA-OV [23] | T | I | zero-shot | 18.6 | 31.2 | 24.8 | 19.5 | 16.7 | 36.3 | 24.0 |
| SplatTalk | T | I | zero-shot | 19.6 | 60.3 | 44.0 | 35.8 | 35.5 | 61.8 | 41.8 |
| SplatTalk-ScanQA-FT | T | I | zero-shot | 28.9 | 66.0 | 43.2 | 28.3 | 33.8 | 60.0 | 41.5 |
| SplatTalk-3DVQA-FT | T | I | zero-shot | 24.3 | 57.0 | 36.1 | 7.8 | 16.4 | 35.0 | 26.5 |
3.3 Entropy Adaptive Gaussian Sampling
SplatTalk leverages FreeSplat [46]’s Gaussian triplet fusion module that significantly reduces the number of 3D Gaussians needed to represent a scene. However, effectively capturing fine details in complex scenes still requires a large number of Gaussians, often exceeding the context window size available for visual tokens in the LLM. Further, not all Gaussians are equally important for downstream 3D-language tasks. To effectively capture the most semantically informative aspects of the scene, we introduce an entropy-based adaptive Gaussian sampling strategy. Specifically, we select the top Gaussians whose language features exhibit the highest entropy, where matches the number of visual tokens accepted by the LLM.
3.4 Optional - Finetuning on 3D VQA Datasets
While SplatTalk is capable of performing the 3D VQA task in a zero-shot manner, it also allows for flexible fine-tuning on 3D VQA datasets to enhance alignment and reasoning capabilities. Rather than training the model end-to-end, which is both computationally expensive and prone to overfitting on limited datasets, we optimize only the LoRA parameters applied to the LLM during fine-tuning.
4 Experiments
4.1 Implementation Details
We train SplatTalk with 500 scenes from the ScanQA training set. The initial self-supervised stage focuses purely on learning the 3D scene representation, without leveraging any text-annotated information. For each scene, we uniformly sample 100 views around the scene and extract their per-patch features by passing them through the visual encoder of LLaVA-OV followed by the multi-modal projector. The subsequent SplatTalk model is trained on a single NVIDIA H100 (80GB) GPU. Both the LoRA fine-tuning and inference are also conducted on a single H100. More details can be found in Supp. Material.
| Sampling Method | EM@1 | EM@1-R | METEOR | ROUGE | CIDEr | SPICE |
| Random | 18.6 | 33.5 | 13.5 | 31.7 | 59.9 | 13.0 |
| Point Density | 17.9 | 34.4 | 14.2 | 32.1 | 60.8 | 14.8 |
| FPS | 18.9 | 34.8 | 14.4 | 33.0 | 62.7 | 14.2 |
| Entropy | 19.8 | 34.0 | 14.6 | 34.4 | 65.1 | 14.7 |
| Visual Length | ScanQA (val) | SQA3D (test) | MSQA | |||
| EM@1 | EM@1-R | EM@1 | EM@1-R | EM@1 | EM@1-R | |
| tokens | 19.8 | 34.0 | 24.5 | 38.2 | 11.9 | 19.1 |
| tokens | 19.8 | 34.0 | 27.8 | 41.2 | 11.9 | 19.1 |
4.2 Datasets and Metrics
Evaluation Dataset. We evaluate our method on ScanNet-based QA datasets: ScanQA [2], SQA3D [34], and MSR3D [29]. ScanQA is a benchmark designed for question answering in 3D indoor scenes, containing detailed spatial and semantic information extracted from ScanNet. SQA3D and MSR3D further extend this by introducing scenarios that describe an embodied agent’s 3D position and state.
Metrics. For ScanQA and SQA3D, we evaluate performance using standard n-gram-based metrics, specifically CIDEr [41], METEOR [3], ROUGE [28], EM@1 [15], and EM@1-Refined for ScanQA, and EM@1 and EM@1-Refined for SQA3D. In contrast, for MSR3D, we utilize the LLM-Match score [31] as recommended by the authors.
4.3 Performance on 3D Question Answering Tasks
ScanQA. We evaluate SplatTalk’s performance against several model families, each representing different levels of specialization and generalization. “Specialist Models” refer to models specifically trained to solve ScanQA without exposure to other tasks. “Generalist 3D LMMs” consist of models trained across a diverse range of 3D vision-language tasks, enabling broader generalization. “Finetuned 3D LMMs” are models that have been further fine-tuned on the ScanQA training set to optimize performance for this particular dataset. Lastly, “2D LMM-Based Models” include models that process only image inputs without leveraging depth or explicit 3D information. SplatTalk and its fine-tuned variants fall within this category.
We evaluate three different variants of our model on this task. The base model (SplatTalk) is trained solely to encode the semantic representation of the scene from multiple RGB images using 3DGS, without exposure to question-answering tasks. Note that this model is trained entirely in a self-supervised manner, without any access to language supervision or textual annotations during training. To further adapt it to downstream 3D VQA tasks, we fine-tune this base model on the ScanQA training set, referring to it as SplatTalk-ScanQA-FT, and on the combined training sets of ScanQA and SQA3D, denoting as SplatTalk-3DVQA-FT. We present results in Table 1.
Compared to LLaVA-OV, whose 2D language feature maps are used to train SplatTalk, our base model, which benefits from a more 3D-aware representation due to the 3D Gaussian training process, achieves significantly higher performance (e.g. vs on CIDEr). Our language fine-tuned model variants, SplatTalk-ScanQA-FT and SplatTalk-3DVQA-FT, improve performance further. Notably, despite not being trained with depth supervision or explicit 3D information, our models achieve comparable results to both Generalist 3D LMMs and Finetuned 3D LMMs, while outperforming Specialist Models. Additionally, among 2D LMM-Based approaches, our base model without being fine-tuned with 3D VQA datasets outperforms other methods, with significantly higher performance on some metrics, including both open-source and closed-source models. We show qualitative results in Fig. 3.
SQA3D. Similar to our evaluation on ScanQA, we compare our model against both 3D LMMs and 2D LMMs. Compared to 2D LLaVA-OV features, our base model with 3D-aware features achieves a significant performance gain, highlighting the advantage of incorporating 3D information via 3DGS which aligns with our findings on ScanQA. Further fine-tuning on ScanQA alone and on the combined ScanQA + SQA3D dataset leads to substantial improvements. Notably, SplatTalk-3DVQA-FT surpasses specialist models that leverage 3D point cloud inputs, demonstrating the strength of our approach in 3D scene understanding.
MSR3D. We evaluated our model’s performance against LLaVA-OV and LEO111The reported number reflects performance across all splits of MSR3D. However, since LEO was trained on ScanNet and 3RScan, its performance remains largely consistent across splits. [19] on the Scannet split of MSR3D. Importantly, all comparisons are conducted in a zero-shot setting, as none of the models have been trained on MSR3D. This is an interleaved dataset where inputs consist of both text and images. However, since all models are tested in a zero-shot manner, they rely solely on text inputs, where the agent’s state descriptions and the question are combined into a single text prompt. On the visual processing side, LEO operates on 3D point cloud inputs, whereas both LLaVA-OV and our model rely only on RGB images.
The results, presented in Table 2, demonstrate that our base model outperforms LLaVA-OV across all question types. Notably, aside from the ”counting” category, our base model achieves approximately twice the performance of LLaVA-OV on other question types. Both models significantly outperform LEO, which struggles due to the absence of fine-tuning on this dataset.
Interestingly, we observe a slight performance drop when comparing SplatTalk with SplatTalk-ScanQA-FT. However, this drop becomes much more pronounced when compared with SlatTalk-3DVQA-FT. A plausible explanation is that fine-tuning on other QA datasets introduces overfitting, causing the model to adapt too specifically to those datasets. Furthermore, the observed performance drop suggests that MSR3D differs significantly from ScanQA and SQA3D, making direct knowledge transfer less effective. This distinction may also explain why LEO struggles to generalize in a zero-shot setting. Despite being trained on multiple QA datasets, none of them closely resemble MSR3D, limiting its ability to adapt without fine-tuning.
4.4 Ablation Studies
Gaussian Sampling Strategies. We explore various Gaussian sampling strategies for selecting tokens as inputs to the LLM, including: (1) Random Sampling—Gaussians are selected random across the scene. (2) Point Density Adaptive Sampling—Gaussians are sampled proportionally to the density of 3D points. (3) Furthest Point Sampling (FPS)—Gaussians are iteratively selected at locations maximally distant from previously chosen points. (4) Entropy Adaptive Sampling—Gaussians are prioritized based on language feature entropy, selecting points with the highest uncertainty.
We conduct experiments on a subset of the ScanQA validation set, consisting of 10 scenes. The results, presented in Table 4, demonstrate that entropy-adaptive sampling consistently outperforms all other sampling strategies. This finding highlights entropy-adaptive sampling as the most effective approach for selecting informative tokens, making it the optimal strategy for downstream 3D VQA tasks.
Length of Visual Tokens. We evaluate the performance of SplatTalk on ScanQA, SQA3D, and MSR3D under two different settings: (1) using only image-worth of visual tokens ( tokens) and (2) using the full context length corresponding to images-worth of visual tokens ( tokens). The results, presented in Table 4, evaluate model performance using EM@1 and EM@1-Refined, as these metrics offer a reliable measure of how precisely model predictions align with ground truth answers, both in literal accuracy and permissible variations in phrasing. We employ entropy-adaptive sampling (see Sec. 3.3), which prioritizes high-information tokens and remains deterministic given a fixed set of visual tokens.
A key observation from our results is that, apart from SQA3D, the performance on ScanQA and MSR3D remains unchanged regardless of the number of tokens used. Specifically, using only image-worth of high-entropy tokens achieves the same EM@1 score as using images-worth of tokens. This suggests that additional scene context does not contribute to performance improvements, which could be due to two reasons: 1) Dataset limitations—These datasets may inherently lack the need for additional scene or spatial information to improve performance; 2) Model limitations—The model may struggle to effectively utilize additional visual tokens for enhanced spatial understanding. In contrast, SQA3D shows a clear performance gain when incorporating more tokens, suggesting that this dataset genuinely requires richer 3D scene context to improve model understanding.
5 Conclusion
In this work, we introduce the first method to integrate free-form language features into a 3D Gaussian Splatting framework, enabling zero-shot generalization for 3D Visual Question Answering (3D VQA) without requiring 3D point clouds or depth maps. Our approach surpasses methods that rely solely on 2D language feature maps and achieves performance on par with 3D LMMs that explicitly utilize 3D structural information across multiple 3D VQA benchmarks. We hope this work inspires further research into leveraging multi-view RGB inputs to encapsulate rich semantic representations of 3D scenes, ultimately enhancing 3D VQA capabilities without the reliance on explicit geometric priors.
Appendix A Data Details
We train SplatTalk on 500 ScanNet scenes from the ScanQA training set. This stage does not involve any language supervisory signal and is purely self-supervised. We evaluate our method and baselines on 3 datasets: ScanQA [2] validation set, SQA3D [34] test set, and MSR3D [29] test set. The ScanQA validation set comprises 71 ScanNet scenes, while the SQA3D and MSR3D test sets includes 67 scenes, with significant overlap with ScanQA validation set.
Appendix B Training & Model Details
SplatTalk Training. The training of SplatTalk consists of the following main steps: 1) Extract high-dimensional language features for each RGB image using LLaVA-OV [23], 2) Train an autoencoder to compress the high-dimensional features obtained from step 1 into a more compact space, and 3) Train the 3D Gaussian Splatting (3DGS) model.
First, to extract language-informed visual features, we leverage LLaVA-OV [23], which processes RGB images through its SigLIP [50]-based vision encoder, followed by a multimodal projector. This transformation ensures that the extracted visual tokens are natively structured for the LLM’s input space, allowing seamless processing and comprehension by the underlying Qwen2 [11] model. These features serve as pseudo 2D ground truth, guiding the training of the 3D-language Gaussian Splatting model.
After obtaining these high-dimensional semantic features, we train an autoencoder to compress them while preserving their most informative components. Each feature map of size , where represents the feature depth, is then flattened, resulting in feature vectors, each of size .
The autoencoder is composed of multiple linear layers, each interleaved with a batch normalization (BN) layer and a GeLU activation function. Specifically,
where indicates the size of the linear layer, BN stands for Batch Norm and GeLU stands for the GeLU activation function. The compressed features are normalized to reside within a unit hypersphere, which we find to be important for stabilizing the 3DGS training process.
We train the autoencoder with batch size , using AdamW [32] optimizer with learning rate for epochs on a single NVIDIA H100 (80GB) GPU.
SplatTalk builds upon the FreeSplat [46] architecture, extending it with a language feature head attached to the Gaussian latent decoder. This additional head predicts a -dimensional feature vector for each Gaussian, alongside the standard Gaussian parameters used for RGB splatting.
For training the 3DGS component, we uniformly sample 100 views per scene. Each input image is resized to resolution to balance efficiency and feature retention. The language features are trained using compressed pseudo 2D ground truth as supervisory signal. Inspired by [51], we implement a parallel CUDA rasterizer pipeline, allowing RGB and language features to be rendered in parallel using the shared Gaussian parameters. This approach is memory efficient and allows for rendering higher-dimensional features than the traditional rendering pipeline.
SplatTalk is trained using the Adam optimizer with an initial learning rate of , which decays smoothly following a cosine decay schedule. We employ a combination of photometric and semantic losses to effectively train SplatTalk. The photometric loss is computed using a combination of MSE and LPIPs, and semantic loss using MSE and cosine distance. The final loss is
where and are the ground-truth RGB image and the pseudo ground-truth language feature map respectively. and correspond to the predicted RGB image and language feature map respectively. We train SplatTalk on a one NVIDIA H100 (80GB) GPU.
SplatTalk Inference. We begin by extracting the reconstructed 3D Gaussians that represent the entire scene. Next, we retrieve the language features corresponding to each 3D Gaussian, capturing the semantic information embedded within the scene representation.
3D VQA Inference. We sample tokens from the set of Gaussian tokens obtained from the SplatTalk inference step. This is equivalent to image tokens, which is the context window size of the LLM. The optimal sampling method we use is entropy-based sampling. We first compute the entropy for each 3D Gaussian. Then we rank them in descending order from highest to lowest entropy. Finally, we sample the top tokens from this list as the input visual tokens to the LLM, along with the language tokens.
3D VQA Fine-tuning. During fine-tuning we employ LoRA [17] with and . The model is fine-tuned for epoch with learning rate on a single NVIDIA H100 (80GB) GPU.
Appendix C Evaluation Metrics
We evaluate model performance on ScanQA using n-gram-based metrics, including CIDEr [41], METEOR [3], ROUGE [28], EM@1 [15], and EM@1-Refined. For SQA3D, we report EM@1 and EM@1-Refined, while for MSR3D, we follow the authors’ recommendation and use the LLM-Match metric [31] to assess answer quality.
Specifically, CIDEr evaluates the generated response by comparing n-gram similarity against multiple reference responses, assigning higher weight to words that frequently appear across multiple references. In contrast, METEOR aligns words between the generated response and reference answers using exact matches, stemming, synonyms, and paraphrasing, providing a more flexible and semantically aware evaluation. It computes both precision and recall, with a higher emphasis on recall, as human evaluation favors covering all key reference words. METEOR captures both semantic and syntactic similarity, making it more robust than purely n-gram-based metrics. ROUGE-L, on the other hand, measures similarity using the Longest Common Subsequence (LCS) between the generated response and the reference answers.
On the other hand, EM@1 or Exact Match at 1, is a strict evaluation metric that only assigns a score if the generated response perfectly matches the reference answer, without any deviations. Even if the response is semantically correct but differs in phrasing, formatting, or minor details, the model is penalized. In contrast, EM@1-Refined offers a more flexible evaluation, allowing for slight variations in wording.
The LLM-Match metric, introduced in [31], uses GPT-4 to assign a score from to to the generated responses based on their similarity to the reference answers. A score of indicates a completely incorrect response, while a score of corresponds to a fully correct answer. Intermediate scores reflect varying degrees of alignment.
Appendix D More Visualization on ScanQA
In Fig. 4 and Fig. 5, we show more qualitative examples on ScanQA. Our SplatTalk can reason about the spatial relationships between objects in the scene significantly better than LLaVA-OV in many cases.